OpenStudy (anonymous):

Identify the factors of x2 − 4x − 12? (x + 4)(x − 3) (x − 4)(x + 3) (x − 2)(x + 6) (x + 2)(x − 6)

OpenStudy (anonymous):

To do this in your head: What are the prime factors of -12? -1*2*2*3 What two things added together could be -4, things that must be pieces of the prime factorization above 6 + -2 = -4 Does 6 * -2 = -12? Yes. 6 = 2*3 -2 = -1*2 So all good. And that's how I do it in my head at a glance.